# Luth: Efficient French Specialization for Small Language Models and Cross-Lingual Transfer

### v1

@@ Abstract
The landscape of Large Language Models (LLMs) remains predominantly English-centric, resulting in a significant performance gap for other major languages, such as French, especially in the context of Small Language Models (SLMs). Existing multilingual models demonstrate considerably lower performance in French compared to English, and research on efficient adaptation methods for French remains limited. To address this, we introduce Luth, a family of French-specialized SLMs: through targeted post-training on curated, high-quality French data, our models outperform all open-source counterparts of comparable size on multiple French benchmarks while retaining their original English capabilities. We further show that strategic model merging enhances performance in both languages, establishing Luth as a new state of the art for French SLMs and a robust baseline for future French-language research.
@@
Maxence Lasbordes Université Paris-Dauphine Télécom SudParis maxence.lasbordes@dauphine.eu          Sinoué Gad École Polytechnique Télécom SudParis sinoue.gad@polytechnique.edu
@@ 1 Introduction
In parallel, Small Language Models have emerged as a promising direction. Studies show that smaller models, when properly trained or adapted, can achieve competitive performance across diverse tasks (Lepagnol et al., 2024; Nguyen et al., 2024). Their compact size enables faster inference, lower computational overhead, and practical deployment, making them well-suited for real-world applications (Belcak et al., 2025). SLMs can also be efficiently specialized to specific languages or domains, offering a practical path to high-quality French language models without relying on large-scale resources.
2 Contributions
In this paper, we introduce Luth [^1] https://github.com/kurakurai/Luth, a family of compact French Small Language Models designed to address the English-centric bias through targeted adaptation. We demonstrate that using carefully curated post-training data, it is possible to significantly improve French capabilities, including general knowledge, instruction-following, and mathematical reasoning, without degrading original English performance, and even enhancing both languages through strategic model merging.
@@ 2 Contributions
    The Luth [^3] https://huggingface.co/collections/kurakurai/luth-models-68d1645498905a2091887a71 family, comprising 5 models ranging from 350M to 1.7B parameters, achieving state-of-the-art performance in French within their size categories and delivering an absolute average improvement of up to $11.26\%$ across six French benchmarks.
@@
    An efficient and reproducible methodology for language-specific adaptation, easily extendable to other mid- and low-resource languages, while preserving performance in other languages.
@@ 3 Related Work
The development of multilingual and language-specific models aims to mitigate the English-centric bias of current LLMs. Recent models such as BLOOM Le Scao et al. (2022), Llama Grattafiori et al. (2024) and AYA Üstün et al. (2024) cover dozens of under-represented languages, but they do not focus on language-specific optimization and frequently underperform on individual languages. Regional initiatives, for example EuroLLM [^4] EuroLLM and CroissantLLM fall within the category of Small Language Models, with parameter counts of 1.7B and 1.3B, respectively., target European languages (including French) to improve coverage and performance for that region Martins et al. (2024). Several efforts concentrate specifically on French. CroissantLLM Faysse et al. (2024) introduced a French–English bilingual model, Lucie Gouvert et al. (2025) has released and open-sourced a large portion of the resources needed for French LLM development, and Pensez Ha (2025) studied French model development with an emphasis on reasoning and argued for quality-over-quantity approaches to data curation. Despite these contributions, important gaps remain. Many works prioritize larger, resource-intensive models or report performance shortfalls relative to multilingual baselines of comparable size. They also provide few practical, low-cost recipes to substantially improve French-language capabilities, leaving room for work on compact, French-specialized models and efficient adaptation strategies suitable for resource-constrained deployments.
@@ 4 Luth-SFT Dataset
To address the lack of high-quality open-source French post-training datasets, we introduce Luth-SFT, comprising 570570k samples ($338$ million tokens) of French instruction–response pairs (Figure 2).
@@
Data Gathering  To construct this dataset, we first collected samples from existing multilingual datasets, including AYA Üstün et al. (2024), Smoltalk2, and CroissantLLM Faysse et al. (2024). These datasets contain texts in multiple languages; to efficiently isolate the French samples among hundreds of thousands of entries, we employed the langdetect library.
@@
Data Translation  To further diversify and expand our French dataset, we selected two high-quality, openly available English instruction datasets, Tülu 3 Lambert et al. (2024) and OpenHermes. Our approach is twofold: (1) translate the English prompts into French (A.1) using strong multilingual models (GPT-4o and Qwen3 32B in non-reasoning mode), and (2) generate new French responses from scratch conditioned on the translated prompts, rather than directly translating the original answers. For Tülu 3, we focused exclusively on the math and instruction-following subsets, as these align with our objectives. The samples produced through this pipeline constitute the majority of our dataset. Notably, for OpenHermes, an existing French version generated with GPT-4o following this methodology was already available, substantially reducing the associated computational cost Alhajar (2025).
@@
Filtering  We used a two-stage filtering pipeline to ensure both dataset quality and domain relevance. The first stage, linguistic validation, enforced strict French language criteria, including grammatical correctness, coherence, absence of code-switching or mixed-language content, and proper instructional formatting. The second stage, content filtering, systematically removed samples from three categories: programming-related content (e.g., code snippets, debugging queries, tool discussions), tool-calling content (e.g., API usage, command-line operations, system configuration), and samples containing logical inconsistencies or factual errors. This approach preserved instruction-following, scientific discourse, and general conversational samples while maintaining high linguistic and content standards. All system prompts used are listed in A.2.
@@
Scholar  This subset was developed to address the scarcity of high-quality scientific resources in French. The dataset draws extensively from Baccalauréat and Classes Préparatoires aux Grandes Écoles (CPGE) examination materials, providing both questions and detailed solutions (see example snippet in A.3) across a broad range of subjects. A key objective was to build a resource that is non-synthetic and rooted in expert knowledge. Examination materials were particularly well-suited for this purpose, as they are typically accompanied by official solutions in PDF format, authored and validated by domain experts. In total, more than $\mathbf{14{,}000}\,\text{PDFs}$ were collected, covering examination sessions from 1980 to 2025555Mainly sourced from Sujet Bac and UPS Sujet.. These documents were processed through a multi-step pipeline (prompts listed in Appendix A):
@@
After processing, the dataset comprises $\mathbf{30{,}300~\text{samples}}$, distributed across several domains. The subject distribution is summarized Table 1. It should be noted that the proportions mainly reflect the availability of data for each subject, and do not represent a deliberate choice on our part.
Table 1: Distribution of scholars by subject.
@@ 5.1 Model Training
As this work focuses on Small Language Models (SLMs) with fewer than 2B parameters, we conducted comprehensive evaluations of multilingual models in this size range to identify the best-performing model for French and to enhance its capabilities. We considered LFM2 (350M, 700M, and 1.2B) LiquidAI (2025) and Qwen3 (0.6B and 1.7B) Yang et al. (2025) for their strong French and English performance. While other SLMs, such as LLaMA 3.2 (1B) Grattafiori et al. (2024), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen2.5 (0.5B and 1.5B) Yang et al. (2024a), are also viable alternatives, our evaluations indicate that they underperform relative to more recent models on the tasks considered in this work. The models were selected based on their capabilities in Math, General Knowledge and Instruction Following in both French and English. Qwen3 and LFM2 variants then went through a full fine-tuning stage, instead of LoRA Hu et al. (2021) for better learning, on our Luth-SFT dataset, which infuses them with a richer understanding of French, specific vocabulary, domain-specific terminology, and improved their skills in the previously mentioned areas.
@@ 6 Evaluation
Model
IFEval
French
GPQA-Diamond
French
MMLU
French
Math500
French
Arc-Challenge
French
Hellaswag
French
@@
To broaden our evaluation beyond French benchmarks, we also focused on English to assess our models’ capabilities. Our evaluation process is fully transparent, and all reported results are reproducible using open-source code [^6] https://github.com/kurakurai/Luth and publicly available data.
@@ 6.1 Benchmark Selection
6.2 Evaluation workflow
@@ 6.3 Results
Model
IFEval
English
GPQA-Diamond
English
MMLU
English
Math500
English
Arc-Challenge
English
Hellaswag
English
@@ 7 Conclusion
This paper introduces Luth, a family of state-of-the-art French Small Language Models (SLMs) that outperform all other models of comparable size on six French benchmarks covering general knowledge, instruction following, and mathematics. Although specialized in French, these models retain strong capabilities in other languages, particularly English, even showing improvements on various English benchmarks through cross-lingual transfer. These results stem from two key innovations: (1) the Luth-SFT, a French post-training dataset which drastically improves the model’s performance in French and (2) the use of model merging to retain multilingual skills while further improving each component’s specialized language capabilities. Moreover, we demonstrate that careful fine-tuning on a specific language alone can yield significant performance gains without resorting to costly methods like continual pretraining. We expect that similar improvements could extend to larger architectures and other languages; verifying this remains a direction for future work.
@@ 8 Limitations
Moreover, we assessed stability primarily in English, without thoroughly evaluating whether the models retain their abilities in other languages. Our experiments were also restricted to Small Language Models (under 2 billion parameters), which may limit the extent to which our approach unlocks potential gains at larger scales.
Finally, the Luth-SFT dataset does not cover key capabilities such as tool use or code generation, which are increasingly central to modern LLMs. While it significantly improves performance in French, it may not fully capture the language’s diverse structures and usage patterns.

### v2

@@ Abstract
The landscape of Large Language Models remains predominantly English-centric, resulting in a significant performance gap for other major languages, such as French, especially in the context of Small Language Models (SLMs). Existing multilingual models demonstrate considerably lower performance in French compared to English, and research on efficient adaptation methods for French remains limited. To address this, we introduce Luth, a family of French-specialized SLMs: through targeted post-training on curated, high-quality French data, our models outperform all open-source counterparts of comparable size on multiple French benchmarks while retaining their original English capabilities. We further show that strategic model merging enhances performance in both languages, establishing Luth as a new state of the art for French SLMs and a robust baseline for future French-language research.
@@
Maxence Lasbordes thanks: Equal contribution. LightOn, Paris Inria Paris maxence.lasbordes@lighton.ai          Sinoué Gad 11footnotemark: 1 École Polytechnique Institut Polytechnique de Paris sinoue.gad@polytechnique.edu
@@ 1 Introduction
In parallel, SLMs have emerged as a promising direction. Studies show that smaller models, when properly trained or adapted, can achieve competitive performance across diverse tasks (Lepagnol et al., 2024; Nguyen and others, 2024). Their compact size enables faster inference, lower computational overhead, and practical deployment, making them well-suited for real-world applications (Belcak et al., 2025). SLMs can also be efficiently specialized to specific languages or domains, offering a practical path to high-quality French language models without relying on large-scale resources.
In this paper, we introduce Luth [^1] https://github.com/kurakurai/Luth, a family of compact French SLMs designed to address the English-centric bias through targeted adaptation. We demonstrate that using carefully curated post-training data, it is possible to significantly improve French capabilities, including general knowledge, instruction-following, and mathematical reasoning, without degrading original English performance, and even enhancing both languages through strategic model merging.
@@
    The Luth [^3] https://huggingface.co/collections/kurakurai/luth-models-68d1645498905a2091887a71 family, including 5 models ranging from 350M to 1.7B parameters, achieving state-of-the-art performance in French within their size categories and delivering an absolute average improvement of up to $11.26\%$ across six French benchmarks.
@@
    An efficient and reproducible methodology for language-specific adaptation, easily extendable to other languages, while preserving performance in other languages.
@@ 2 Related Work
The development of multilingual and language-specific models aims to mitigate the English-centric bias of current LLMs. Models such as BLOOM Le Scao et al. (2022), Llama Grattafiori et al. (2024), and AYA Üstün et al. (2024) cover dozens of under-represented languages, but they do not focus on language-specific optimization and often underperform on individual languages. Regional initiatives, such as EuroLLM Martins et al. (2024) and Apertus Apertus et al. (2025), aim to improve multilingual coverage, with Apertus supporting over 1,000 languages and emphasizing data compliance.
Several efforts focus specifically on French. Early work includes PAGnol Launay et al. (2021), which introduced scaling laws for French and trained a 1.5B-parameter GPT model. More recent contributions include CroissantLLM Faysse et al. (2024), a French–English bilingual model; Gaperon Godey et al. (2025), a fully open suite of French–English–code models emphasizing transparency and reproducibility; Lucie Gouvert et al. (2025), which open-sourced substantial resources for French LLM development; and Pensez Ha (2025), which studied French models with a focus on reasoning and data quality.
Despite these contributions, important gaps remain. Many works prioritize large, resource-intensive models or report performance shortfalls relative to multilingual baselines of comparable size. Moreover, they offer few practical, low-cost recipes to substantially improve French-language capabilities, leaving room for compact, French-specialized models and efficient adaptation strategies suitable for resource-constrained settings.
@@ 3 Luth-SFT Dataset
To address the lack of high-quality open-source French post-training datasets, we introduce Luth-SFT, which contains 570570k samples ($338$ million tokens) of French instruction–response pairs (Figure 2).
@@
Data Gathering  To build this dataset, we first collected parts from existing multilingual datasets, including AYA Üstün et al. (2024), Smoltalk2 HuggingFaceTB (2025), and CroissantLLM Faysse et al. (2024). As the datasets are massively multilingual, we language filtered the French samples via the langdetect libary Danilak (2021).
@@
Data Translation  To further diversify and expand our French dataset, we selected two high-quality, openly available English instruction datasets, Tülu 3 Lambert et al. (2024) and OpenHermes Teknium (2023). Our approach is twofold: (1) translate the English prompts into French (A.1) using strong multilingual models (GPT-4o and Qwen3 32B in non-reasoning mode), and (2) generate new French responses from scratch conditioned on the translated prompts, rather than directly translating the original answers. For Tülu 3, we focused exclusively on the math and instruction-following subsets, as these align with our objectives. The samples produced through this pipeline constitute the majority of our dataset. Notably, for OpenHermes, an existing French version generated with GPT-4o following this methodology was already available, substantially reducing the associated computational cost Alhajar (2025).
@@
Filtering  We used a two-stage filtering pipeline to ensure both dataset quality and domain relevance. The first stage, linguistic validation, enforces strict French language criteria, including grammatical correctness, coherence, absence of code-switching or mixed-language content, and proper instructional formatting. The second stage, content filtering, systematically removes samples from three categories: programming-related content (e.g., code snippets, debugging queries, tool discussions), tool-calling content (e.g., API usage, command-line operations, system configuration), and samples containing logical inconsistencies or factual errors. This approach preserved instruction-following, scientific discourse, and general conversational samples while maintaining high linguistic and content standards. All system prompts used are listed in A.2.
@@
Scholar  This subset was developed to address the scarcity of high-quality scientific resources in French. The dataset draws extensively from Baccalauréat and Classes Préparatoires aux Grandes Écoles (CPGE) examination materials, providing both questions and detailed solutions (see example snippet in A.3) across a broad range of subjects. A key objective was to build a resource that is non-synthetic and rooted in expert knowledge. Examination materials were particularly well-suited for this purpose, as they are typically accompanied by official solutions in PDF format, authored and validated by domain experts. In total, more than $\mathbf{14{,}000}\,\text{PDFs}$ were collected, covering examination sessions from 1980 to 2025444Mainly sourced from Sujet Bac and UPS Sujet.. These documents were processed through a multi-step pipeline (prompts listed in Appendix A):
@@
After processing, the dataset contains $\mathbf{30{,}300~\text{samples}}$, distributed across several domains. The subject distribution is summarized Table 1. It should be noted that the proportions mainly reflect the availability of data for each subject, and do not represent a deliberate choice on our part.
@@ 4.1 Model Training
As this work focuses on SLMs with fewer than 2B parameters, we conducted comprehensive evaluations of multilingual models in this size range to identify the best-performing model for French and to enhance its capabilities. We considered LFM2 (350M, 700M, and 1.2B) LiquidAI (2025) and Qwen3 (0.6B and 1.7B) Yang et al. (2025) for their strong French and English performance. While other SLMs, such as LLaMA 3.2 (1B) Grattafiori et al. (2024), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen2.5 (0.5B and 1.5B) Yang et al. (2024a), are also viable alternatives, our evaluations indicate that they underperform relative to more recent models on the tasks considered in this work. The models were selected based on their capabilities in Math, General Knowledge and Instruction Following in both French and English. Qwen3 and LFM2 variants then went through a full fine-tuning stage, instead of LoRA Hu et al. (2021) for better learning, on our Luth-SFT dataset, which infuses them with a richer understanding of French, specific vocabulary, domain-specific terminology, and improved their skills in the previously mentioned areas.
@@ 5 Evaluation
Model IFEval French GPQA-Diamond French MMLU French Math500 French Arc-Challenge French Hellaswag French
@@
As the models we test were trained on a large part of English data, we also evaluate on English to assess our model’s capabillities on that language after having been optimized in French with our techniques. Our evaluation process is fully transparent, and all reported results are reproducible using open-source code [^5] https://github.com/kurakurai/Luth and publicly available data.
@@ 5.1 Benchmark Selection
5.2 Evaluation workflow and Reasonning mode
@@ 5.3 Results
Model IFEval English GPQA-Diamond English MMLU English Math500 English Arc-Challenge English Hellaswag English
@@ 6 Conclusion
This paper introduces Luth, a family of state-of-the-art French SLMs that outperform all other models of comparable size on six French benchmarks covering general knowledge, instruction following, and mathematics. Although specialized in French, these models retain strong capabilities in other languages, particularly English, even showing improvements on various English benchmarks through cross-lingual transfer. These results stem from two key innovations: (1) the Luth-SFT, a French post-training dataset which drastically improves the model’s performance in French and (2) the use of model merging to retain multilingual skills while further improving each component’s specialized language capabilities. Moreover, we demonstrate that careful fine-tuning on a specific language alone can yield significant performance gains without resorting to costly methods like continual pretraining. We expect that similar improvements could extend to larger architectures and other languages; verifying this remains a direction for future work.
@@ 7 Limitations
Moreover, we assessed stability primarily in English without thoroughly evaluating whether the models retain their ability in other languages. Our experiments were also restricted to SLMs (under 2 billion parameters), which may limit the extent to which our approach unlocks potential gains at larger scales.
Finally, the Luth-SFT dataset does not cover key capabilities such as tool use or code generation, which are increasingly central to modern LLMs.
Acknowledgments
We thank Djamé Seddah and Thibaud Southiratn for comments on earlier versions of this work. This work was partly funded by the BPI Scribe project.